Solve for A B C:
4a+7b=1
8a-2c=-38
6b-2c=20
How do I begin to do this?

Answer by ewatrrr(24785)


Hi
solving sytem by Substitution
4a+7b=1 4a = 1-7b
8a-2c=-38 c = 4a +19
6b-2c=20 c = 3b -10
4a + 19 = 3b - 10
(1-7b) + 19 = 3b - 10
30 = 10b
b = 3
a = -5 | a = (1-7b)/4
c = -1 | c = 3b -10